# Analysis and Predictive Modeling of Solar Coronal Holes Using Computer Vision and LSTM Networks

### v2

@@ Abstract
In the era of space exploration, coronal holes on the sun play a significant role due to their impact on satellites and aircraft through their open magnetic fields and increased solar wind emissions. This study employs computer vision techniques to detect coronal hole regions and estimate their sizes using imagery from the Solar Dynamics Observatory (SDO). Additionally, we utilize deep learning methods, specifically Long Short-Term Memory (LSTM) networks, to analyze trends in the area of coronal holes and predict their areas across various solar regions over a span of seven days. By examining time series data, we aim to identify patterns in coronal hole behavior and understand their potential effects on space weather. This research enhances our ability to anticipate and prepare for space weather events that could affect Earth’s technological systems.
@@ 1 Introduction
Addressing this gap, our study combines computer vision with deep learning to predict coronal hole sizes. Automated detection ensures accuracy and consistency, eliminating manual inspection errors, and enhancing efficiency. Using the Long Short-Term Memory (LSTM) model (Hochreiter and Schmidhuber 1997), we forecast coronal hole sizes over a week, providing insights into impending space weather events. Our scalable method allows for extensive dataset analysis, offering insights into long-term coronal hole patterns. This research highlights the critical role of AI in enhancing predictive capabilities for space weather events, ultimately protecting Earth’s technological infrastructure (Oceanic and Administration ).
@@ 3.1 Coronal Hole Detection using Computer Vision
In our segmentation, the middle region, which includes the Left Area, Redzone, and Right Area, generally involves areas close to the solar equator, potentially extending to about ±30° latitude from the solar equator as shown in Figure 2. These regions are more likely to impact Earth’s geomagnetic environment directly due to their equatorial proximity. Specifically, the Redzone is defined within a 15° wide longitudinal slice around the heliocentric meridian and falls within ±30° latitude from the solar equator (Nakagawa, Nozawa, and Shinbori 2019). Areas outside of this middle region are classified as polar regions. Additionally, for quantitative analysis, we calculate the percentage of the white area (coronal holes) relative to the total solar surface area, considering the entire solar area as a unity 1.
@@
Polar Region. The polar regions of the sun also exhibit distinct behaviors, as shown in Figure 6. The ”polarity rule” demonstrates that coronal hole areas at the north and south poles often behave oppositely, aligning with their magnetic field configurations and affecting solar wind distribution.
These findings emphasize the importance of continuous monitoring and further research to understand the long-term impacts of coronal holes on space weather and geomagnetic disturbances on Earth
@@ 3.2 Time Series Prediction using Deep Learning
In our study, we complement computer vision techniques with deep learning to predict the future behavior of coronal holes using LSTM networks (Hochreiter and Schmidhuber 1997). LSTMs efficiently manage long-term dependencies in this data, crucial for accurate predictions (Yang et al. 2019; Torres, Martínez-Álvarez, and Troncoso 2022). The architecture of LSTM involves several gates—input, forget, and output—that control how information is stored, updated, or discarded, making it highly effective for our sequential data on coronal holes (Hochreiter and Schmidhuber 1997).
@@ 4 Results
We employ Long Short-Term Memory (LSTM) (Hochreiter and Schmidhuber 1997) networks to predict the area of coronal holes in the sun’s middle region. The LSTM model, trained on 3857 days of numerical data from the middle regions of AIA 193 observations extracted by the BCH model, is adept at handling time-series data, capturing patterns and trends in solar activity.
Figure 7: Comparison of actual and predicted middle area values for a 7-day period from July 31, 2021, to August 06, 2021.
Figure 8: Comparison of actual and predicted middle area values for a 7-day period from July 31, 2021, to August 06, 2021, obtained after training the LSTM on data from January 06, 2011, to July 30, 2021.
Figure 7 and Figure 8 show the LSTM model’s accuracy over a week-long period, aligning closely with actual observations. The predictive modeling of coronal hole areas is not only academically interesting but also practically important. The size and location of coronal holes can influence the occurrence and trajectory of CMEs (Webb and Howard 2012) and CIRs (Heber, Sanderson, and Zhang 1999), which can have significant effects on Earth’s space weather, including geomagnetic storms and auroras in the future.
@@ 5 Discussion
This study enhances understanding of coronal holes and their impacts on space weather by developing a detection and prediction model. While effective, the model focuses mainly on mid-zone areas, neglecting variations in regions like the polar regions which might influence space weather. Additionally, it covers only one solar cycle, limiting its scope. Future research should delve deeper into the physical processes governing coronal hole formation, particularly the sun’s magnetic dynamics. Expanding the dataset to cover multiple solar cycles and employing more sophisticated forecasting tools would enhance predictive capabilities, allowing for better preparation against space weather events impacting Earth.
@@ 6 Conclusion
A simplified model for detecting coronal holes is essential for enhancing space weather forecasting. Coronal holes, marked by their low temperature and density, influence solar winds that can affect Earth’s satellites and power grids. Accurate detection and measurement of coronal holes are crucial for predicting space weather events and mitigating their impacts. Our research developed a model that identifies coronal holes, generating 3857 data entries. This data was used to create a deep learning prediction model, specifically using LSTM networks, to estimate the size of coronal holes. The model showed promising results, effectively predicting the coronal hole size using numerical data. This model improves our ability to predict space weather events, aiding in the prevention of potential impacts on Earth’s technological infrastructure. Additionally, understanding the size and location of coronal holes helps in predicting solar activities like CMEs and CIRs.

### v3

# Analysis and Predictive Modeling of Solar Coronal Holes Using Computer Vision and ARIMA-LSTM Networks

@@ Abstract
In the era of space exploration, coronal holes on the sun play a significant role due to their impact on satellites and aircraft through their open magnetic fields and increased solar wind emissions. This study employs computer vision techniques to detect coronal hole regions and estimate their sizes using imagery from the Solar Dynamics Observatory (SDO). Additionally, we utilize hybrid time series prediction model, specifically combination of Long Short-Term Memory (LSTM) networks and ARIMA, to analyze trends in the area of coronal holes and predict their areas across various solar regions over a span of seven days. By examining time series data, we aim to identify patterns in coronal hole behavior and understand their potential effects on space weather.
@@ 1 Introduction
Addressing this gap, our study combines computer vision with deep learning to predict coronal hole sizes. Automated detection ensures accuracy and consistency, eliminating manual inspection errors, and enhancing efficiency. Using the Long Short-Term Memory (LSTM) model (Hochreiter and Schmidhuber 1997) and ARIAM (Box and Jenkins 1970), we forecast coronal hole sizes over a week, providing insights into impending space weather events. Our scalable method allows for extensive dataset analysis, offering insights into long-term coronal hole patterns. This research highlights the critical role of AI in enhancing predictive capabilities for space weather events, ultimately protecting Earth’s technological infrastructure (Oceanic and Administration ).
@@ 3.1 Coronal Hole Detection using Computer Vision
In our segmentation, the middle region, which includes the Left Area, Redzone, and Right Area, generally involves areas close to the solar equator, potentially extending to about ±30° latitude from the solar equator as shown in Figure 2. These regions are more likely to impact Earth’s geomagnetic environment directly due to their equatorial proximity.
Specifically, the Redzone is defined within a 15° wide longitudinal slice around the heliocentric meridian and falls within ±30° latitude from the solar equator (Nakagawa, Nozawa, and Shinbori 2019). Areas outside of this middle region are classified as polar regions. Additionally, for quantitative analysis, we calculate the percentage of the white area (coronal holes) relative to the total solar surface area, considering the entire solar area as a unity 1.
@@
Polar Region. The polar regions of the sun also exhibit distinct behaviors, as shown in Figure 6. The ”polarity rule” demonstrates that coronal hole areas at the north and south poles often behave oppositely, aligning with their magnetic field configurations and affecting solar wind distribution. These findings emphasize the importance of continuous monitoring and further research to understand the long-term impacts of coronal holes on space weather and geomagnetic disturbances on Earth
Figure 7: The plots from A to B show the ARIMA-LSTM predicted the next 7 days of middle area of coronal holes. Each plot corresponds to a different time period used for training: (A) 2011-01-06 to 2019-04-11, (B) 2011-01-06 to 2020-02-02. The blue lines indicate the actual coronal hole areas used for training and the actual areas for the following 7 days. The red lines show the predicted areas for the next 7 days based on the given training data. The green lines represent the data used for training. This visualization highlights the effectiveness of the ARIMA-LSTM hybrid model in predicting the future behavior of coronal holes.
@@ 3.2 Time Series Prediction using Deep Learning
In our study, we integrate Long Short-Term Memory (LSTM) networks (Hochreiter and Schmidhuber 1997) with AutoRegressive Integrated Moving Average (ARIMA) (Box and Jenkins 1970) models to predict the future behavior of coronal holes. LSTM networks efficiently manage long-term dependencies in data, crucial for accurate time series predictions (Yang et al. 2019; Torres, Martínez-Álvarez, and Troncoso 2022). Additionally, we use the pm.auto_arima function from the pmdarima library (Smith et al. 2019) to automatically determine the best fitting parameters for the ARIMA model. For our predictions, we generate forecasts from both the ARIMA and LSTM models over a 7-day period and then compute the average of these predictions to obtain the final forecasted values. This hybrid approach enhances the overall prediction accuracy by leveraging the strengths of both models.
@@ 4 Results
4.1 Experimental Setting
In our study, we employed a Long Short-Term Memory (LSTM) network to predict the future behavior of coronal holes. The LSTM model architecture consists of an initial LSTM layer with 50 units and a ReLU activation function, configured to return sequences. This is followed by a Dropout layer with a dropout rate of 0.2 to prevent overfitting. Next, there is another LSTM layer with 50 units and a ReLU activation function, which does not return sequences. Another Dropout layer with a dropout rate of 0.2 is added, and the final layer is a Dense layer with a single unit for the output. The model is compiled using the mean squared error (MSE) loss function and the Adam optimizer, with the learning rate set to the default value. The model was trained for 50 epochs with a batch size of 128 and a learning rate of 0.01.
4.2 Experimental Results
We employ ARIMA-LSTM (Hochreiter and Schmidhuber 1997; Box and Jenkins 1970) networks to predict the area of coronal holes in the sun’s middle region. The LSTM model, trained on 3857 days of numerical data from the middle regions of AIA 193 observations extracted by the BCH model, is adept at handling time-series data, capturing patterns and trends in solar activity.
Figure 8: The plots from C to D show the ARIMA-LSTM predicted the next 7 days of middle area of coronal holes. Each plot corresponds to a different time period used for training: (C) 2011-01-06 to 2020-08-29, (D) 2011-01-06 to 2021-03-06. The blue lines indicate the actual coronal hole areas used for training and the actual areas for the following 7 days. The red lines show the predicted areas for the next 7 days based on the given training data. The green lines represent the data used for training. This visualization highlights the effectiveness of the ARIMA-LSTM hybrid model in predicting the future behavior of coronal holes.
Figure 7, 8 illustrates the performance of the ARIMA-LSTM hybrid model in predicting the Daily middle area of coronal holes for the next 7 days. The model is trained on different time periods, as shown in plots A to D, which cover various spans from 2011-01-06 to 2021-03-06. The blue lines represent the actual coronal hole areas used for training and the actual areas for the subsequent 7 days. The red lines indicate the predicted areas for the next 7 days based on the training data, and the green lines represent the data used for training.
The patterns observed in the coronal hole areas suggest that these features exhibit specific trends and behaviors over time. By leveraging these patterns, the ARIMA-LSTM hybrid model can effectively learn and predict future coronal hole areas. This approach demonstrates that it is possible to accurately predict the behavior of coronal holes based on their historical patterns, providing valuable insights for understanding and forecasting space weather phenomena.
@@ 5 Discussion
This study enhances understanding of coronal holes and their impacts on space weather by developing a detection and prediction model. While effective, the model focuses mainly on mid-zone areas, neglecting variations in regions like the polar regions which might influence space weather. Additionally, it covers only one solar cycle, limiting its scope. Future research should delve deeper into the physical processes governing coronal hole formation, particularly the sun’s magnetic dynamics.
@@ 6 Conclusion
A simplified model for detecting coronal holes is essential for enhancing space weather forecasting. Coronal holes, marked by their low temperature and density, influence solar winds that can affect Earth’s satellites and power grids. Accurate detection and measurement of coronal holes are crucial for predicting space weather events and mitigating their impacts. Our research developed a model that identifies coronal holes, generating 3857 data entries. This data was used to create a deep learning prediction model, specifically using ARIMA-LSTM, to estimate the size of coronal holes. The model showed promising results, effectively predicting the coronal hole size using numerical data. This model improves our ability to predict space weather events, aiding in the prevention of potential impacts on Earth’s technological infrastructure. Additionally, understanding the size and location of coronal holes helps in predicting solar activities like CMEs and CIRs.